# allinspirekeys.bib

### v1

@@ 4 Results
Figure 10: Combined continuum limits for the plaquette and $\beta$ for every combination of extrapolation set and small volume extrapolation. Further explanations in the text.
@@ 5 Discussion
Still, there might be a systematic effect unaccounted for, because all our extrapolation result lie below the Hamiltonian curve. One possible explanation for this systematic deviation could be the truncation on the Hamiltonian side. This seems not to be the case because larger $l$-values tend to push the plaquette values up at fixed $\beta$. Despite the discussion from above, we certainly cannot be 100% sure that we are close enough to the temporal continuum limit, which might offer one explanation for the systematic deviation. However, this will need further investigation in the future.
@@ 6 Summary and Outlook
We have performed the temporal continuum limit in a U$(1)$ lattice gauge theory using stochastic simulations in the Lagrangian formalism. We performed this temporal continuum limit using the anisotropic lattice formulation starting with two $\beta$-values from the isotropic side. Trajectories of constant spatial lattice spacing are defined by keeping the Sommer parameter $r_{0}/a_{s}$ fixed. The so obtained temporal continuum results for the plaquette and the coupling $\beta$ are compared to results from a direct Hamiltonian simulation. We find general agreement within two $\sigma$ between Hamiltonian and extrapolated Lagrangian results. As discussed in the previous section, the deviation is systematic towards lower plaquette values for the extrapolated results, for which we currently do not have a good explanation.
@@ Appendix A Autocorrelation
We take care of the autocorrelation in a different way: The potential energies are still determined with bootstrapping, but with block length 1. At the same time, we determine the autocorrelation of the original data with the UWerr-algorithm [Wolff:2003sm] implemented in [kostrzewa_2024_14800686]. We then rescale the results of the bootstrapping to take the autocorrelation into account. The error is rescaled with $2\tau_{int}$ and each bootstrap sample $x$ is rescaled to be $x_{\mathrm{new}}=x_{\mathrm{old}}+(x_{\mathrm{old}}-\mu)\cdot(2\cdot\tau_{int}-1)$, with $\mu$ the unbiased mean. This ensures that the difference of the bootstrap sample to the unbiased mean is $2\tau_{int}$ times as large, so the error increases by the required amount.
We also draw bootstrap samples of the plaquette. There, we set the block length to $4\tau_{int}^{2}$ to take the autocorrelation into account.
@@ Appendix B Using AIC to determine the potential points
The potential points are computed with the help of the Akaike Information Criterion (AIC) [Borsanyi:2020mff].
@@ C.1 Getting a touching ellipse
We know the Lagrangian result $(\beta,P)$ and its errors $(\sigma_{\beta},\sigma_{P})$. This is the centre point of the ellipse, and the ratio of the errors is the ratio of the major axes of the ellipse. The angle $\phi$ of the ellipsis is given by $\tan(2\phi)=\frac{2\rho\sigma_{\beta}\sigma_{P}}{\sigma_{\beta}^{2}-\sigma_{P}^{2}}$, with $\rho$ the correlation coefficient of the bootstrap samples of $\beta$ and $P$ [CowanGlen1998SDA].
@@
A polar to a point $P$ can be written as $P^{T}A\tilde{X}=0$ [Fortuna:2016].
@@
the pole has the coordinates $(\frac{x_{P}}{z_{P}},\frac{y_{P}}{z_{P}})$, where we chose $z_{P}$ as our rescaling factor for the polar equation. If the polar is a tangent to the ellipse, the pole is the touching point and directly on the ellipse [Fortuna:2016]. In that case, it fulfils the ellipse equation eq. 24.
@@ Appendix D Systematics of continuum limit
$\xi_{\text{input}}$ $\beta$ confs. $L$ $T$ algo # sweeps $\tau_{\text{int}}$
$1$ $1.65$ $5500$ $16$ $16$ met $50$ $0.490(13)$
$1$ $1.7$ $9748$ $16$ $16$ met $100$ $0.497(10)$
$4/5$ $1.59$ $1500$ $16$ $20$ met $50$ $0.516(65)$
$4/5$ $1.6$ $1500$ $16$ $20$ met $50$ $0.585(95)$
$4/5$ $1.615$ $1500$ $16$ $20$ met $50$ $0.499(45)$
$4/5$ $1.63$ $1500$ $16$ $20$ met $50$ $0.489(44)$
$4/5$ $1.64$ $9749$ $16$ $20$ met $100$ $0.490(10)$
$4/5$ $1.65$ $4749$ $16$ $20$ met $100$ $0.502(39)$
$4/5$ $1.6521$ $1750$ $16$ $20$ met $50$ $0.496(34)$
$4/5$ $1.665$ $1500$ $16$ $20$ met $50$ $0.497(63)$
$4/5$ $1.7$ $4749$ $16$ $20$ met $100$ $0.503(39)$
$4/5$ $1.75$ $4749$ $16$ $20$ met $100$ $0.500(25)$
@@
$\xi_{\text{input}}$ $\beta$ confs. $L$ $T$ algo # sweeps $\tau_{\text{int}}$
$1$ $1.65$ $9667$ $3$ $16$ met $50$ $0.4818(99)$
$1$ $1.7$ $9667$ $3$ $16$ met $50$ $0.497(10)$
$4/5$ $1.6$ $7501$ $3$ $20$ met $100$ $0.491(20)$
$4/5$ $1.615$ $7501$ $3$ $20$ met $100$ $0.489(23)$
$4/5$ $1.64$ $7501$ $3$ $20$ met $100$ $0.522(36)$
$4/5$ $1.65$ $9667$ $3$ $20$ met $50$ $0.500(20)$
$4/5$ $1.6521$ $7501$ $3$ $20$ met $100$ $0.490(11)$
$2/3$ $1.56$ $7251$ $3$ $24$ met $100$ $0.481(11)$
$2/3$ $1.5614$ $7251$ $3$ $24$ met $100$ $0.502(33)$
$2/3$ $1.565$ $7251$ $3$ $24$ met $100$ $0.496(31)$
$2/3$ $1.58$ $9167$ $3$ $24$ met $50$ $0.501(23)$
$2/3$ $1.595$ $46250$ $3$ $24$ hb $20$ $0.4944(80)$
$2/3$ $1.6$ $7251$ $3$ $24$ met $100$ $0.498(31)$
$2/3$ $1.6075$ $7251$ $3$ $24$ met $100$ $0.499(12)$
$1/2$ $1.49$ $9667$ $3$ $32$ met $100$ $0.496(10)$
$1/2$ $1.515$ $9667$ $3$ $32$ met $100$ $0.493(10)$
$1/2$ $1.5381$ $9667$ $3$ $32$ met $100$ $0.607(44)$
$1/2$ $1.54$ $8667$ $3$ $32$ met $50$ $0.493(21)$
$1/2$ $1.55$ $8667$ $3$ $32$ met $50$ $0.498(11)$
$1/2$ $1.5514$ $9667$ $3$ $32$ met $100$ $0.532(32)$
$1/2$ $1.5531$ $8667$ $3$ $32$ met $50$ $0.634(51)$
$1/2$ $1.555$ $9667$ $3$ $32$ met $100$ $0.491(10)$
$2/5$ $1.46$ $14501$ $3$ $40$ met $100$ $0.502(24)$
$2/5$ $1.4638$ $14501$ $3$ $40$ met $100$ $0.724(51)$
$2/5$ $1.47$ $14501$ $3$ $40$ met $100$ $0.507(22)$
$2/5$ $1.48$ $14501$ $3$ $40$ met $100$ $0.667(45)$
$2/5$ $1.51$ $7667$ $3$ $40$ met $50$ $0.573(43)$
$2/5$ $1.511$ $7667$ $3$ $40$ met $50$ $0.524(34)$
$2/5$ $1.52$ $14501$ $3$ $40$ met $100$ $0.539(28)$
$2/5$ $1.525$ $14501$ $3$ $40$ met $100$ $0.729(51)$
$1/3$ $1.44$ $37001$ $3$ $48$ met $100$ $0.605(24)$
$1/3$ $1.4406$ $37001$ $3$ $48$ met $100$ $1.407(93)$
$1/3$ $1.45$ $37001$ $3$ $48$ met $100$ $1.310(83)$
$1/3$ $1.46$ $4667$ $3$ $48$ met $50$ $0.791(95)$
$1/3$ $1.48$ $4667$ $3$ $48$ met $50$ $0.93(12)$
$1/3$ $1.4814$ $37001$ $3$ $48$ met $100$ $1.326(85)$
$1/3$ $1.5$ $37001$ $3$ $48$ met $100$ $1.292(82)$
$1/4$ $1.42$ $49334$ $3$ $64$ met $100$ $9.6(13)$
$1/4$ $1.43$ $29336$ $3$ $64$ met $50$ $4.53(55)$
$1/4$ $1.4378$ $2667$ $3$ $64$ met $50$ $2.39(61)$
$1/4$ $1.47$ $49334$ $3$ $64$ met $100$ $1.72(11)$
$1/4$ $1.49$ $49334$ $3$ $64$ met $100$ $12.8(19)$
$1/5$ $1.41$ $47499$ $3$ $80$ hb $200$ $0.820(37)$
$1/5$ $1.42$ $47500$ $3$ $80$ hb $200$ $0.868(40)$
$1/5$ $1.46$ $47499$ $3$ $80$ hb $200$ $0.923(44)$
$1/5$ $1.47$ $47500$ $3$ $80$ hb $200$ $0.964(47)$
$1/5$ $1.48$ $47500$ $3$ $80$ hb $200$ $0.986(48)$
$0.18$ $1.39$ $36770$ $3$ $88$ hb $500$ $1.349(87)$
$0.18$ $1.395$ $36000$ $3$ $88$ hb $500$ $1.284(81)$
$0.18$ $1.4$ $36000$ $3$ $88$ hb $500$ $1.420(95)$
$0.18$ $1.405$ $19742$ $3$ $88$ hb $500$ $1.28(11)$
$0.18$ $1.45$ $36756$ $3$ $88$ hb $500$ $1.69(12)$
$0.18$ $1.46$ $37277$ $3$ $88$ hb $500$ $1.99(15)$
$0.18$ $1.47$ $36000$ $3$ $88$ hb $500$ $1.75(13)$
[…]

### v2

@@ 4 Results
Figure 10: Combined continuum limits for the plaquette and $\beta$ for every combination of extrapolation set and small volume extrapolation. Further explanations are in the text.
@@ 5 Discussion
Still, there might be a systematic effect unaccounted for, because all our extrapolation results lie below the Hamiltonian curve. One possible explanation for this systematic deviation could be the truncation on the Hamiltonian side. This seems not to be the case because larger $l$-values tend to push the plaquette values up at fixed $\beta$. Despite the discussion from above, we certainly cannot be 100% sure that we are close enough to the temporal continuum limit, which might offer one explanation for the systematic deviation. However, this will need further investigation in the future.
@@ 6 Summary and Outlook
We have performed the temporal continuum limit in a U(1)(1) lattice gauge theory using stochastic simulations in the Lagrangian formalism. We performed this temporal continuum limit using the anisotropic lattice formulation starting with two $\beta$-values from the isotropic side. Trajectories of constant spatial lattice spacing are defined by keeping the Sommer parameter $r_{0}/a_{s}$ fixed. The so obtained temporal continuum results for the plaquette and the coupling $\beta$ are compared to results from a direct Hamiltonian simulation. We find general agreement within two $\sigma$ between Hamiltonian and extrapolated Lagrangian results. As discussed in the previous section, the deviation is systematic towards lower plaquette values for the extrapolated results, for which we currently do not have a good explanation. Our procedure allows us to calculate observables at the same lattice spacing $a_{s}$ in both theories. For the comparison between the theories, we are not interested in the limit $a_{s}\to 0$, but instead we use the matching point to combine advantages of both theories at fixed lattice spacing.
@@ Acknowledgements
This version of the article has been accepted for publication, after peer review but is not the Version of Record and does not reflect post-acceptance improvements, or any corrections. The Version of Record is available online at: http://dx.doi.org/10.1140/epjc/s10052-025-14923-2.
@@ Appendix A Autocorrelation
We take care of the autocorrelation in a different way: The potential energies are still determined with bootstrapping, but with block length 1. At the same time, we determine the autocorrelation of the original data with the UWerr-algorithm [40] implemented in [39]. We then rescale the results of the bootstrapping to take the autocorrelation into account. The error is rescaled with $2\tau_{\text{int}}$ and each bootstrap sample $x$ is rescaled to be $x_{\mathrm{new}}=x_{\mathrm{old}}+(x_{\mathrm{old}}-\mu)\cdot(2\cdot\tau_{\text{int}}-1)$, with $\mu$ the unbiased mean. This ensures that the difference of the bootstrap sample to the unbiased mean is $2\tau_{\text{int}}$ times as large, so the error increases by the required amount.
We also draw bootstrap samples of the plaquette. There, we set the block length to $4\tau_{\text{int}}^{2}$ to take the autocorrelation into account.
@@ Appendix B Using AIC to determine the potential points
The potential points are computed with the help of the Akaike Information Criterion (AIC) [41].
@@ C.1 Getting a touching ellipse
We know the Lagrangian result $(\beta,P)$ and its errors $(\sigma_{\beta},\sigma_{P})$. This is the centre point of the ellipse, and the ratio of the errors is the ratio of the major axes of the ellipse. The angle $\phi$ of the ellipsis is given by $\tan(2\phi)=\frac{2\rho\sigma_{\beta}\sigma_{P}}{\sigma_{\beta}^{2}-\sigma_{P}^{2}}$, with $\rho$ the correlation coefficient of the bootstrap samples of $\beta$ and $P$ [42].
@@
A polar to a point $P$ can be written as $P^{T}A\tilde{X}=0$ [43].
@@
the pole has the coordinates $(\frac{x_{P}}{z_{P}},\frac{y_{P}}{z_{P}})$, where we chose $z_{P}$ as our rescaling factor for the polar equation. If the polar is a tangent to the ellipse, the pole is the touching point and directly on the ellipse [43]. In that case, it fulfils the ellipse equation eq. 24.
@@ Appendix D Systematics of continuum limit
$\xi_{\text{input}}$ $\beta$ # confs. $L$ $T$ algo # sweeps $\tau_{\text{int}}$
11 $1.65$ $5500$ $16$ $16$ met $50$ $0.490(13)$
11 $1.7$ $9748$ $16$ $16$ met $100$ $0.497(10)$
$4/5$ $1.59$ $1500$ $16$ 2020 met $50$ $0.516(65)$
$4/5$ $1.6$ $1500$ $16$ 2020 met $50$ $0.585(95)$
$4/5$ $1.615$ $1500$ $16$ 2020 met $50$ $0.499(45)$
$4/5$ $1.63$ $1500$ $16$ 2020 met $50$ $0.489(44)$
$4/5$ $1.64$ $9749$ $16$ 2020 met $100$ $0.490(10)$
$4/5$ $1.65$ $4749$ $16$ 2020 met $100$ $0.502(39)$
$4/5$ $1.6521$ $1750$ $16$ 2020 met $50$ $0.496(34)$
$4/5$ $1.665$ $1500$ $16$ 2020 met $50$ $0.497(63)$
$4/5$ $1.7$ $4749$ $16$ 2020 met $100$ $0.503(39)$
$4/5$ $1.75$ $4749$ $16$ 2020 met $100$ $0.500(25)$
@@
$\xi_{\text{input}}$ $\beta$ # confs. $L$ $T$ algo # sweeps $\tau_{\text{int}}$
11 $1.65$ $9667$ 33 $16$ met $50$ $0.4818(99)$
11 $1.7$ $9667$ 33 $16$ met $50$ $0.497(10)$
$4/5$ $1.6$ $7501$ 33 2020 met $100$ $0.491(20)$
$4/5$ $1.615$ $7501$ 33 2020 met $100$ $0.489(23)$
$4/5$ $1.64$ $7501$ 33 2020 met $100$ $0.522(36)$
$4/5$ $1.65$ $9667$ 33 2020 met $50$ $0.500(20)$
$4/5$ $1.6521$ $7501$ 33 2020 met $100$ $0.490(11)$
$2/3$ $1.56$ $7251$ 33 $24$ met $100$ $0.481(11)$
$2/3$ $1.5614$ $7251$ 33 $24$ met $100$ $0.502(33)$
$2/3$ $1.565$ $7251$ 33 $24$ met $100$ $0.496(31)$
$2/3$ $1.58$ $9167$ 33 $24$ met $50$ $0.501(23)$
$2/3$ $1.595$ $46250$ 33 $24$ hb 2020 $0.4944(80)$
$2/3$ $1.6$ $7251$ 33 $24$ met $100$ $0.498(31)$
$2/3$ $1.6075$ $7251$ 33 $24$ met $100$ $0.499(12)$
$1/2$ $1.49$ $9667$ 33 $32$ met $100$ $0.496(10)$
$1/2$ $1.515$ $9667$ 33 $32$ met $100$ $0.493(10)$
$1/2$ $1.5381$ $9667$ 33 $32$ met $100$ $0.607(44)$
$1/2$ $1.54$ $8667$ 33 $32$ met $50$ $0.493(21)$
$1/2$ $1.55$ $8667$ 33 $32$ met $50$ $0.498(11)$
$1/2$ $1.5514$ $9667$ 33 $32$ met $100$ $0.532(32)$
$1/2$ $1.5531$ $8667$ 33 $32$ met $50$ $0.634(51)$
$1/2$ $1.555$ $9667$ 33 $32$ met $100$ $0.491(10)$
$2/5$ $1.46$ $14501$ 33 $40$ met $100$ $0.502(24)$
$2/5$ $1.4638$ $14501$ 33 $40$ met $100$ $0.724(51)$
$2/5$ $1.47$ $14501$ 33 $40$ met $100$ $0.507(22)$
$2/5$ $1.48$ $14501$ 33 $40$ met $100$ $0.667(45)$
$2/5$ $1.51$ $7667$ 33 $40$ met $50$ $0.573(43)$
$2/5$ $1.511$ $7667$ 33 $40$ met $50$ $0.524(34)$
$2/5$ $1.52$ $14501$ 33 $40$ met $100$ $0.539(28)$
$2/5$ $1.525$ $14501$ 33 $40$ met $100$ $0.729(51)$
$1/3$ $1.44$ $37001$ 33 $48$ met $100$ $0.605(24)$
$1/3$ $1.4406$ $37001$ 33 $48$ met $100$ $1.407(93)$
$1/3$ $1.45$ $37001$ 33 $48$ met $100$ $1.310(83)$
$1/3$ $1.46$ $4667$ 33 $48$ met $50$ $0.791(95)$
$1/3$ $1.48$ $4667$ 33 $48$ met $50$ $0.93(12)$
$1/3$ $1.4814$ $37001$ 33 $48$ met $100$ $1.326(85)$
$1/3$ $1.5$ $37001$ 33 $48$ met $100$ $1.292(82)$
$1/4$ $1.42$ $49334$ 33 $64$ met $100$ $9.6(13)$
$1/4$ $1.43$ $29336$ 33 $64$ met $50$ $4.53(55)$
$1/4$ $1.4378$ $2667$ 33 $64$ met $50$ $2.39(61)$
$1/4$ $1.47$ $49334$ 33 $64$ met $100$ $1.72(11)$
$1/4$ $1.49$ $49334$ 33 $64$ met $100$ $12.8(19)$
$1/5$ $1.41$ $47499$ 33 $80$ hb $200$ $0.820(37)$
$1/5$ $1.42$ $47500$ 33 $80$ hb $200$ $0.868(40)$
$1/5$ $1.46$ $47499$ 33 $80$ hb $200$ $0.923(44)$
$1/5$ $1.47$ $47500$ 33 $80$ hb $200$ $0.964(47)$
$1/5$ $1.48$ $47500$ 33 $80$ hb $200$ $0.986(48)$
$0.18$ $1.39$ $36770$ 33 $88$ hb $500$ $1.349(87)$
$0.18$ $1.395$ $36000$ 33 $88$ hb $500$ $1.284(81)$
$0.18$ $1.4$ $36000$ 33 $88$ hb $500$ $1.420(95)$
$0.18$ $1.405$ $19742$ 33 $88$ hb $500$ $1.28(11)$
$0.18$ $1.45$ $36756$ 33 $88$ hb $500$ $1.69(12)$
$0.18$ $1.46$ $37277$ 33 $88$ hb $500$ $1.99(15)$
$0.18$ $1.47$ $36000$ 33 $88$ hb $500$ $1.75(13)$